OpenStudy (anonymous):

f(x)=1/x−5 and g(x)=x+(2/x). what is g of f? isnt it...1/(x-5) + 2/1/(x-5) ???

OpenStudy (anonymous):

\[\frac{1}{x-5} + (\frac{2}{\frac{1}{x-5}})\]

OpenStudy (anonymous):

So you're correct :)

OpenStudy (anonymous):

friggin software is not accepting my answer

OpenStudy (phi):

you should simplify that ugly second fraction...

OpenStudy (anonymous):

You might have to simplify that mess though :/

OpenStudy (anonymous):

i see

OpenStudy (phi):

multiply top and bottom by (x-5)

OpenStudy (anonymous):

thanks